OpenStudy (anonymous):

I can't solve this trig equation, can somebody help. sec(x)-tan(x)=sqr3/3

OpenStudy (anonymous):

\[\sec(x)-\tan(x)=\frac{ \sqrt{3} }{ 3 }\]

OpenStudy (anonymous):

@IMStuck

OpenStudy (imstuck):

give me one sec ok? I will work on it while I help this other person get through his dilemmma then I will be here, ok?

OpenStudy (anonymous):

Yeah no problem

OpenStudy (imstuck):

i'm here...working on it now...

OpenStudy (imstuck):

what exactly are you doing with this? Proving that the left side equals the right side?

OpenStudy (anonymous):

no the book says to solve the trigonometry equation.

OpenStudy (anonymous):

using trig identities

OpenStudy (imstuck):

I am not finding any way in which to get this to a quadratic type of equation to solve. I get this when I use the identities:\[\frac{ 1-\sin(x) }{ \cos(x) }=\frac{ \sqrt{3} }{ 3 }\]And you need to get the sin and the cos to be either one trig function or the other. There is no identity to use that combines a single sin and a single cos into one trig function. Not so far as I know, at least. Are you sure you didn't omit a ^2 in there or something?

OpenStudy (anonymous):

No, i have tried too and i don't know to what trig identity to relate it to.

OpenStudy (anonymous):

In one of the examples of the book it says to make it equal to zero.